some of the cylindrical silicon wafers used in the semiconductor industry are 10.00 cm in diameter and contain 3.42 × 10^-3 mol si. if the density of silicon is 2.33 g/cm^3, how thick is one of these wafers?
Answer #1

Answer :

Density of silicon = 2.33 g/cm​​​​​3

diameter of silicon wafer = 10 cm

So, radius of silicon wafer = 5 cm

Moles of Si used = 3.42 * 10-3 mole

Molar mass of Si = 28 g/mol

Mass of Si wafers = moles of Si * molar mass = 3.42 * 10-3 * 28 = 95.76 * 10-3 g

As we know : density = mass / volume

So, volume = mass of Si wafer / density of Si =

= 95.76 * 10-3 / 2.33 = 41.1 * 10-3 cm3 ​​​

So, volume of Si wafer = 41.1 * 10-3 cm​​​​​3

It is given that Si wafer are cylinderical in shape and has radius = 5 cm

Let h be thickness of wafer

Volume of Cylinder = π * r​​​​​​2 * h

So, volume of Si wafer = volume of Cylinder

=> 41.1 * 10-3 = π * 52 * h

h = 41.1 * 10-3 / (π * 25) = 5.23 * 10-4 cm

So, thickness of Si wafer = 5.23 * 10-4 cm = 5.23 um

(um = micrometer)